OpenStudy (anonymous):

is 0 – (–9) 9 or -9?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

0 - (-9) is the same as saying 0+9 so 0+9=+9

OpenStudy (anonymous):

think of it as 0 - 1(-9) - * - = + so -1 * -9 = 9 answer i 9

OpenStudy (anonymous):

To my Calculations as 0 - 1(-9) - * - = + so -1 * -9 = 9 answer i 9 just as kimmy rep said :D

OpenStudy (anonymous):

heh I didn't want to explain like that because relating that to multiplying by (-1) may confuse him/her at this math level :)